xsquared +4x+1=6
solve with two linear factors by factoring

Answer by jim_thompson5910(35256)

x^2+4x+1=6

x^2+4x+1-6=0

x^2+4x-5=0

(x-1)(x+5)=0

x-1=0 or x+5=0

x=1 or x=-5